OpenStudy (anonymous):

Factor completely: 2x2 + 9x + 4 (2x + 2)(x + 2) (2x + 1)(x + 4) (2x + 4)(x + 1) (2x + 2)(x + 4)

mathslover (mathslover):

\[\large{2x^2+9x+4=2x^2+8x+x+4}\]

mathslover (mathslover):

right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

ok so can u try to factor that out?

mathslover (mathslover):

\[\large{2x^2+8x+x+4}\] \[\large{2x(x+4)+1(x+4)}\] right?

OpenStudy (anonymous):

yes

mathslover (mathslover):

further factorization ? can u .. basic funda about that: \[\large{a(b+c)+d(b+c)=(a+d)(b+c)}\] can u do that now?

OpenStudy (anonymous):

yes thanks

mathslover (mathslover):

can u tell me what r u getting?

OpenStudy (anonymous):

(2x + 1)(x + 4)

mathslover (mathslover):

very good .. u r right